# Near to Mid-term Risks and Opportunities of Open Source Generative AI

### v1

@@ B.1.2 Biden’s Executive Order
President Biden’s 2023 Executive Order (EO) on Safe, Secure, and Trustworthy Artificial Intelligence (House, 2023) continues to follow a “soft law” approach of earlier EOs, largely trading enforceable regulation for voluntary industry commitments (PricewaterhouseCoopers, 2024). Safety and security measures surrounding AI technology include requirements for developers to share red-teaming results with the US federal government, and for companies working on “dual-use” foundation models (i.e., systems with civilian and military applications) and/or with large compute clusters to provide regular activity reports. The National Institute of Standards and Technology is set up to play a key role in developing standards for secure and safe AI. Instead of placing hard restrictions on the use of certain AI technology (as the EU AI Act explicitly does), Biden’s EO focuses on promoting best practices, evaluations, and standard development across a wide variety of aspects including security and risk mitigation. For example, it includes references to biological weapons, AI-generated content watermarking, and labor market impacts, and, additionally, measures for attracting foreign national AI talent through streamlining visa procedures and by providing assistance to small businesses and developers. National security interests are also formulated, including the reporting of foreign users of US Infratructure as a Service (IaaS) products, as well as promoting the development of AI-driven tools to detect cyber vulnerabilities.
@@ B.1.3 China’s GenAI Legislation
The earliest legal framework specifically targeting GenAI models and systems, the Chinese government’s Provisional Administrative Measures of Generative Artificial Intelligence Services (Generative AI Measures) (Cyberspace Administration of China, ; Translate, 2023), came into force in China in August 2023. These regulations pose strict obligations on providers of GenAI, ranging from outcome-driven provisions (e.g., requiring generative AI services to not produce illegal or untruthful content) to provenance obligations on training data and model weights, and measures targeted to protect intellectual property and privacy rights (Steven Chong, 2023). From the point of view of open-source model developers, the inability to predict future downstream use of models and systems provided introduces legal risks that require regulatory containment. However, perhaps surprisingly given the prominent role of open-source GenAI in the Chinese economy, these regulations do not govern open-source (GP)AI models specifically (Asia Society, 2024).

### v2

@@ B.1.2 Biden’s Executive Order
President Biden’s 2023 Executive Order (EO) on Safe, Secure, and Trustworthy Artificial Intelligence (House, 2023) continues to follow a “soft law” approach of earlier EOs, largely trading enforceable regulation for voluntary industry commitments (PricewaterhouseCoopers, 2024). Safety and security measures surrounding AI technology include requirements for developers to share red-teaming results with the US federal government, and for companies working on “dual-use” foundation models (i.e., systems with civilian and military applications) and/or with large compute clusters to provide regular activity reports. The National Institute of Standards and Technology (NIST) is set up to play a key role in developing standards for secure and safe AI. Instead of placing hard restrictions on the use of certain AI technology (as the EU AI Act explicitly does), Biden’s EO focuses on promoting best practices, evaluations, and standard development across a wide variety of aspects including security and risk mitigation. For example, it includes references to biological weapons, AI-generated content watermarking, and labor market impacts, and, additionally, measures for attracting foreign national AI talent through streamlining visa procedures and by providing assistance to small businesses and developers. National security interests are also formulated, including the reporting of foreign users of US Infrastructure as a Service (IaaS) products, as well as promoting the development of AI-driven tools to detect cyber vulnerabilities.
@@ B.1.3 China’s Gen AI Legislation
The earliest legal framework specifically targeting Gen AI models and systems, the Chinese government’s Provisional Administrative Measures of Generative Artificial Intelligence Services (Generative AI Measures) (Cyberspace Administration of China, ; Translate, 2023), came into force in China in August 2023. These regulations pose strict obligations on providers of Gen AI, ranging from outcome-driven provisions (e.g., requiring generative AI services to not produce illegal or untruthful content) to provenance obligations on training data and model weights, and measures targeted to protect intellectual property and privacy rights (Steven Chong, 2023). From the point of view of open-source model developers, the inability to predict future downstream use of models and systems provided introduces legal risks that require regulatory containment. Although open-source Gen AI plays a significant role in the Chinese economy, however, these regulations do not seem to target open-source (GP)AI models specifically (Asia Society, 2024).